OpenStudy (anonymous):

Five tickets to a baseball game cost $60. How much do 3 tickets cost? A. $12 B. $24 C. $36 D. $48

OpenStudy (anonymous):

oh jkjk

OpenStudy (anonymous):

\[\frac{$60}{5} = $12\]So it costs $12 per person.\[$12\times3 = $36\]

OpenStudy (futureescape):

60/5 = 12 12*3 = 36 C. 36

OpenStudy (anonymous):

60 / 5 = 12 cost 0f 1 12* 3 = ???

OpenStudy (anonymous):

divide 5 by 60 so 60/5 then multiply that answer by 3...that will give u what u need